# DreamMix: Decoupling Object Attributes for Enhanced Editability in Customized Image Inpainting

### v1

@@ Abstract
Subject-driven image inpainting has emerged as a popular task in image editing alongside recent advancements in diffusion models. Previous methods primarily focus on identity preservation but struggle to maintain the editability of inserted objects. In response, this paper introduces DreamMix, a diffusion-based generative model adept at inserting target objects into given scenes at user-specified locations while concurrently enabling arbitrary text-driven modifications to their attributes. In particular, we leverage advanced foundational inpainting models and introduce a disentangled local-global inpainting framework to balance precise local object insertion with effective global visual coherence. Additionally, we propose an Attribute Decoupling Mechanism (ADM) and a Textual Attribute Substitution (TAS) module to improve the diversity and discriminative capability of the text-based attribute guidance, respectively. Extensive experiments demonstrate that DreamMix effectively balances identity preservation and attribute editability across various application scenarios, including object insertion, attribute editing, and small object inpainting. Our code is publicly available at https://github.com/mycfhs/DreamMix.
@@
00footnotetext: Equal contribution.00footnotetext: Corresponding author.
1 Introduction
Recent advancements in diffusion models [14, 13, 34, 30, 31, 28] have catalyzed remarkable progress in image generation, thereby stimulating the evolution in image editing [11, 3, 27, 36]. Among these applications, subject-driven image customization [33, 17, 40, 10] has garnered significant interest, which aims to generate realistic images of a reference object in various contexts. To achieve this, existing methods utilize different adaption strategies to enable pre-trained text-to-image diffusion models to apply various editing effects while maintaining the object’s identity. Despite their efficacy, these methods often focus on regenerating entire scenes, posing challenges when users want to insert specific objects into designated areas within a given scene. This task, known as Subject-driven Image Inpainting, has significant demand in various practical applications, such as effect-image rendering, poster design, virtual try-on, etc.
To address this task, existing methods [39, 6, 5] typically follow the image inpainting protocol, utilizing a source image and a binary mask as inputs. Additionally, exemplar information is extracted from reference images and employed to condition pretrained diffusion models [31, 28]. However, these reference visual features often encompass overly specific content, thereby constraining the generative capacity of pre-trained text-to-image models and complicating the effective editing of objects with desired effects or attributes. Recent efforts [26, 18] have been made to overcome this limitation by incorporating both image and text guidance through attention mechanisms. Nevertheless, challenges remain in balancing training data efficiency with visual editing quality.
@@ 1 Introduction
To overcome the outlined limitations, we propose DreamMix, a diffusion-based generative model for text-subject-driven image inpainting that achieves attribute-level precision and data efficiency. In DreamMix, we introduce a disentangled inpainting framework that separates the process into two stages: Local Content Generation (LCG) and Global Context Harmonization (GCH), enhancing both local subject integration and overall visual coherence. Additionally, we develop an innovative Attribute Decoupling Mechanism (ADM) to diversify the textual description and image pairs of decoupled attributes for the subject, effectively mitigating overfitting issues. At the testing phase, we implement a Textual Attribute Substitution (TAS) module that employs an orthogonal decomposition strategy to separate interfering information from textual guidance, further improving object editing quality. As shown in Fig. 1, the proposed DreamMix demonstrates robust performance across various subject-driven inpainting applications, including identity preservation, attribute editing, and small-object inpainting.
Our main contributions can be summarized as follows:
  • We propose DreamMix, a method that leverages text and image guidance to allow precise object insertion and attribute-level modification within any scene.
  • We introduce an attribute decoupling mechanism and a textual attribute substitution module to mitigate identity overfitting and support diverse attribute editing.
  • Our method demonstrates superior performance over previous approaches across various application scenarios, as evidenced by both quantitative metrics and subjective evaluations.
2 Related Work
Image Customization. Recent advancements in generative models have significantly enhanced image customization techniques, allowing for more nuanced and flexible control over the content of generated images [41, 40, 36, 11, 37, 35]. Recently, subject-driven image customization methods [10, 33, 9, 17] have garnered increasing attention, which utilize diffusion models to produce high-quality and personalized images based on a couple of template images. For instance, Textual Inversion [10] addresses the personalization challenge by using a set of images depicting the same subject to learn a special text token. DreamBooth [33] binds a unique identifier to the concept by finetuning the entire U-Net [32] and the text encoder [29] together with a collection of subject images. However, these methods primarily focus on regenerating entire images but struggle with situations where users want to edit specific local regions of an existing image.
Image Inpainting. Image inpainting involves reconstructing missing or modified regions of an image. Powered by text-to-image diffusion models, recent works [23, 1, 2, 43, 38, 22, 42, 21] have shown significant improvements in repainting the local regions of existing images with text guidance. For example, SD-Inpainting [31] expands the pretrained diffusion models with a concatenation of noise, masked image, and mask as input and retrains it to produce visually consistent inpainting results. PowerPaint [43] enhances the model’s ability to support various inpainting tasks by designing learnable task prompts for text-guided object restoration, context-aware image restoration, and object removal, etc. However, despite their effectiveness in text-driven image inpainting, these methods often encounter limitations in compositing user-specific objects into the desired regions of a scene.
Subject-driven Inpainting. Recent methods have started focusing on combining subject-driven image customization with image inpainting to form a task named subject-driven inpainting [6, 18, 26, 39, 5]. To address this task, some methods [39, 6, 5] utilize an image-guided framework that retrains diffusion models using subject images as templates. However, the absence of original text guidance limits these methods to modify object attributes effectively. This motivates some approaches to develop an image-text-guided architecture that integrates these elements into latent attention layers [26, 18]. Nevertheless, most methods rely on a large-scale training protocol that involves large-scale text-image data construction and a lengthy retraining process. While TIGIC [18] is faster, its tuning-free protocol often struggles to provide consistent editing results. In this paper, we propose DreamMix, which introduces a disentangled inpainting framework and an attribute decoupling mechanism to achieve data-efficient, editing-enhanced subject-driven image inpainting.
3 Method
Given a background image $\boldsymbol{x}$, a binary mask $\boldsymbol{m}$, and a text prompt $\boldsymbol{p}$, DreamMix aims to inpaint the local region of $\boldsymbol{x}$ specified by $\boldsymbol{m}$ to produce an image $\hat{\boldsymbol{x}}$. The refilled areas should not only contain the subject in the reference image(s) $\boldsymbol{x}_{s}$ but also align with the descriptions provided in the text prompt. The overall pipeline of DreamMix is illustrated in Fig. 2. It contains three key modules: a disentangled inpainting framework, an attribute decoupling mechanism, and a textual attribute substitution module, to achieve realistic editable subject-driven inpainting. We describe each module in detail in the following sections.
3.1 Preliminaries
Text-to-Image Diffusion Models. Diffusion models [28, 31] generate samples from random noise by learning to estimate the distribution in the reverse process. Specifically, given a random noise $\boldsymbol{\epsilon}\sim\mathcal{N}({0},\mathbf{I})$ and text prompt $\boldsymbol{p}$, text-to-image diffusion models use a U-Net $\boldsymbol{\epsilon}_{\theta}$ to produce a denoised image. Given a dataset of $\mathcal{D}=\{\boldsymbol{x},\boldsymbol{p}\}$, the entire model is trained using the following objective:
$\mathcal{L_{DM}}\left(\theta;\boldsymbol{x},\boldsymbol{p}\right)=\mathbb{E}_{\boldsymbol{\epsilon}\sim\mathcal{N}({0},\mathbf{I}),t\sim\mathcal{U}({0},\mathbf{T})}\left[||\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t};\boldsymbol{p},t)-\boldsymbol{\epsilon}||_{2}^{2}\right],$ (1)
@@ 3.1 Preliminaries
3.2 Disentangled Inpainting Framework
Existing subject-driven inpainting methods [39, 6, 26] usually build on pretrained diffusion models, wherein the lowest resolution of latent features is 32 times smaller than the original input images. This substantial reduction in resolution hinders these methods’ ability to effectively inpaint objects in vert confined local regions. Additionally, in complex scenes, background interference can affect the editing control of text instructions, leading to imprecise synthesis of the fill-in areas. To overcome these challenges, we propose an effective disentangled inpainting framework, consisting of Local Content Generation (LCG) and Global Context Harmonization (GCH), to enhance both local subject integration and global visual coherence. The overall framework is illustrated in Fig. 2 (c).
Local Content Generation. Some recent research [27, 7, 4] suggests that diffusion models exhibit a potential for course-to-fine generation, characterized by early-stage layout construction, intermediate-stage structure enhancement, and late-stage detail refinement. This insight motivates us to partition the entire inpainting process into two consecutive stages: LCG and GCH. Specifically, we implement LCG at the preceding time steps of $\lambda{T}$, where $\lambda\in[0,1]$ is a trade-off for time step separation. To improve the generation ability across varying-sized areas, we crop the background image $\boldsymbol{x}$ into a local patch and encode it by:
$\boldsymbol{z}^{L}=\text{Enc}(\text{Crop}(\boldsymbol{x}_{s},\boldsymbol{m}^{\prime})).$ (3)
Here, $\text{Crop}(\cdot)$ denotes an image cropping operation guided by a mask $\boldsymbol{m}^{\prime}$, which is produced by slightly enlarging the original local region of $\boldsymbol{m}$ to involve more context. $\text{Enc}(\cdot)$ is the VAE encoder, and $\boldsymbol{z}^{L}$ is the latent code of local image. At each time step, the U-Net $\boldsymbol{\epsilon}_{\theta}$ generates the denoised latent code $\boldsymbol{z}_{t}$ by taking the text prompt as guidance:
$\boldsymbol{z}_{t}=\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z}_{t+1};\boldsymbol{p}_{dec},t),$ (4)
where $\boldsymbol{z}_{t+1}$ is the output at $t+1$ step, and $\boldsymbol{p}_{dec}$ represents the decomposed text embedding that will be introduced in Sec. 3.4. To maintain the input format of pretrained inpainting models, we exploit a blending strategy during inference to ensure that non-editing regions remain unchanged. This process can be formulated as:
where $\odot$ denotes element-wise multiplication and $\boldsymbol{z}_{t}^{L}$ is the noisy latent code of $\boldsymbol{x}^{L}$. $\hat{\boldsymbol{z}}_{t}$ denotes the blended latent code that will be fed to the U-Net in the next time step as in Eq. 4. After the denoising process of $\lambda T$ steps, we project the final result as an RGB image $\hat{\boldsymbol{x}}^{L}$ and repaste it into the original image $\boldsymbol{x}$ to form an intermediate image $\boldsymbol{x}^{G}$, which will be used as an image guidance in the global context harmonization stage.
Global Context Harmonization. The local content generation in the preceding time steps effectively yields accurate object compositing with a well-defined layout and object appearance. However, without incorporating the global image content, it tends to produce a disharmonious “copy-paste” effect in the inpainting areas. To address this issue, we implement a global context harmonization module on the remaining $(1-\lambda)T$ steps to improve the overall harmony between the generated local region and the background image. Similar to LCG, we start by transforming the intermediate image $\boldsymbol{x}^{G}$ into a latent code $\boldsymbol{z}^{G}$ using Eq. 3 as $\boldsymbol{z}^{G}=\text{Enc}(\text{Crop}(\boldsymbol{x}^{G},\boldsymbol{m}))$. Here, $\boldsymbol{m}$ is the original user-specific mask. Then, we adopt the same blending strategy in GCH as follows:
@@ 3.3 Attribute Decoupling Mechanism
Few-shot personalization techniques have gained significant attention for their ability to balance data efficiency with generalization capability. However, limitations remain within the few-shot training protocol. On one hand, the textual descriptions used during finetuning are often overly simplified, providing only basic information about the subject’s appearance. For example, in DreamBooth [33], an image is typically described as “a [classname]”. On the other hand, these methods compress all the object’s details (e.g., color, texture, and shape) into a single identity token [sks]. However, finetuning diffusion models on just a few image-text pairs can lead to subject overfitting, ultimately limiting the models’ generative ability in text-driven editing.
@@
In response, we propose an Attribute Decoupling Mechanism (ADM) to address the data scarcity and subject overfitting issues. To circumvent the heavy workload of manual annotation, we leverage advanced Vision-Language Models [24] to establish an automatic data reconstruction pipeline. The overall framework of ADM is illustrated in Fig. 3. Specifically, given a set of subject image-text pairs $\mathcal{D}_{s}=\{\boldsymbol{x}_{s},\boldsymbol{p}_{s}\}$, we feed the subject images $\boldsymbol{x}_{s}$ into a VLM to generate an attribute dictionary that enumerates all attribute words associated with the given subject. Then, we ask the VLM to randomly combine the detected attributes with the original subject to form a series of text prompts with detailed descriptions such as “a [sks] [attributes] [class name] [attributes]”. Finally, we generate new images based on the reconstructed text prompts to form a regular benchmark $\mathcal{D}_{r}=\{\boldsymbol{x}_{r},\boldsymbol{p}_{r}\}$.
Figure 3: Pipeline of Attribute Decoupling Mechanism (ADM). We obtain the attribute word list using a VLM agent [24] and create regular data with more diverse text formats and image contents.
@@
Combining source data with regular data to finetune the pretrained inpainting models can help reduce overfitting, yet is non-trivial. DreamBooth finetunes the model parameters beyond just identity tokens, which can lead to attribute words mixing up details. For example, in “a brown clay teapot”, the color word might inadvertently bind some texture details, making it difficult to change the texture when transitioning to “a brown glass teapot”, or influence the precise of texture features when switching to “a red clay teapot”. We refer to this phenomenon as “Concept Infusion.” Similarly, SID [16] notes that adding detailed descriptions can disrupt the object’s identity, aligning with this Concept Infusion issue.
@@
To mitigate this issue, we utilize the regular data $\mathcal{D}_{r}$ from ADM for model finetuning and redesign the data reconstruction loss to fit the inpainting task. Specifically, we employ user masks to enforce the inpainting models to fill specific regions. And we propose a loss weight reallocation strategy on Eq. 1 to reduce the influence of background on feature learning:
$\begin{split}\mathcal{L}_{RE}(\theta;\boldsymbol{x},\boldsymbol{p},\boldsymbol{m})=\tau_{1}(\boldsymbol{m}\odot||\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x};\boldsymbol{p})-\boldsymbol{\epsilon}||_{2}^{2})+\\ \tau_{2}((1-\boldsymbol{m})\odot||\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x};\boldsymbol{p})-\boldsymbol{\epsilon}||_{2}^{2}),\end{split}$ (7)
@@
where $\tau_{1}=1.5$ and $\tau_{2}=0.7$ to emphasize the fill-in regions while suppressing the background area. $\hat{\boldsymbol{x}}=\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x};\boldsymbol{p})$ represents the pretrained inpainting model that generates the denosied output $\hat{\boldsymbol{x}}$ from a background image $\boldsymbol{x}$, a text prompt $\boldsymbol{p}$, and a binary mask $\boldsymbol{m}$. As illustrated in Fig. 2 (a), we combine both subject data $\mathcal{D}_{s}$ and regular data $\mathcal{D}_{r}$ with this reweight loss to form an overall loss:

### v2

@@ Abstract
Subject-driven image inpainting has recently gained prominence in image editing with the rapid advancement of diffusion models. Beyond image guidance, recent studies have explored incorporating text guidance to achieve identity-preserved yet locally editable object inpainting. However, these methods still suffer from identity overfitting, where original attributes remain entangled with target textual instructions. To overcome this limitation, we propose DreamMix, a diffusion-based framework adept at inserting target objects into user-specified regions while concurrently enabling arbitrary text-driven attribute modifications. DreamMix introduces three key components: (i) an Attribute Decoupling Mechanism (ADM) that synthesizes diverse attribute-augmented image-text pairs to mitigate overfitting; (ii) a Textual Attribute Substitution (TAS) module that isolates target attributes via orthogonal decomposition, and (iii) a Disentangled Inpainting Framework (DIF) that seperates local generation from global harmonization. Extensive experiments across multiple inpainting backbones demonstrate that DreamMix achieves a superior balance between identity preservation and attribute editability across diverse applications, including object insertion, attribute editing, and small object inpainting.
keywords:
Image editing, Image customization, Diffusion model, Image inpainting.
1 Introduction
Recent advances in diffusion models [10, 12, 46, 41, 40, 35] have catalyzed remarkable progress in image generation, thereby stimulating the evolution in image editing [11, 3, 36, 50, 49, 59, 9, 16]. Among these applications, subject-driven image customization [44, 19, 61, 8, 54] has garnered significant interest, which aims to generate realistic images of a reference object in various contexts. To achieve this, existing methods utilize different adaption strategies to enable pre-trained text-to-image diffusion models to apply various editing effects while maintaining the object’s identity. Despite their efficacy, these methods often focus on regenerating entire scenes, posing challenges when users want to insert specific objects into designated areas within a given scene. This task, known as Subject-driven Image Inpainting, has significant demand in various practical applications, such as effect-image rendering, poster design, virtual try-on, etc.
To address this task, existing methods [58, 5, 4] follow the image inpainting protocol, utilizing a source image and a binary mask as inputs. Additionally, exemplar information is extracted from reference images and used to condition pretrained diffusion models [40, 35]. However, these reference visual features often encompass overly specific content, thereby constraining the generative capacity of pretrained text-to-image models and complicating the effective editing of objects with desired effects or attributes. Recent efforts [37, 26] have been made to overcome this limitation by incorporating image and text guidance through attention mechanisms. Nevertheless, challenges remain in balancing training data efficiency with visual editing quality.
Figure 1: Comparison with state-of-the-art methods in terms of identity preservation, attribute editing, and training data. The reported score is an average on CLIP-T, CLIP-I, and DINO in two applications, respectively. The circle diameter indicates the training data size. Best viewed in color.
In this work, we draw inspiration from recent foundational inpainting models [6, 24], which have demonstrated an impressive ability to fill in image regions with high-quality, semantically consistent content based on textual instructions. Leveraging the power of these pretrained text-driven inpainting models for subject-driven inpainting can help alleviate the data-accuracy trade-off, yet is non-trivial. A potential solution is to adopt parameter-efficient image customization techniques (e.g., DreamBooth [44]) to adapt models with a few template images. However, this straightforward approach encounters certain limitations in achieving precise object insertion and modification. First, the denoising process in inpainting models may introduce global noise that interferes with local information integration when synthesizing target areas. Second, customization strategies like few-shot fine-tuning [44, 8] usually inject the subject’s appearance into a single identity token $\texttt{sks}]$. However, coupling all object attributes into $\texttt{sks}]$ may lead to overfitting on the object’s appearance, ultimately diminishing the effectiveness of text-based editing instructions.
@@ 1 Introduction
To address these limitations, we propose DreamMix, a diffusion-based generative model for text-subject-driven image inpainting that achieves attribute-level precision and data efficiency. In DreamMix, we introduce two key innovations to resolve overfitting and editing conflicts. First, the Attribute Decoupling Mechanism (ADM) addresses data scarcity and overfitting by leveraging Vision-Language Models (VLMs) to generate attribute-diverse training pairs. By reconstructing prompts and synthesizing corresponding regularization images, ADM disentangles entangled attributes during training, preserving the model’s capacity for flexible editing. Second, the Textual Attribute Substitution (TAS) module enhances inference-time control. Through orthogonal decomposition of text embeddings, TAS isolates and suppresses residual attribute signals from the original object, ensuring that editing instructions dominate the generation process. Additionally, to enhance the subject detail and context harmony in arbitrary inpainting regions, we introduce a Disentangled Inpainting Framework (DIF) on the inpainting backbones, which separates the denoising process into local content generation and global context harmonization.
As shown in Figure 1 and 2, DreamMix demonstrates superior performance over state-of-the-art methods in both identity preservation and attribute editing, extending editing ability beyond only ID-preserved methods [4, 5], while consuming significantly less training data than large-scale retraining approaches [37, 30]. In addition, the three proposed components are architecture-agnostic, achieving consistent improvement in various inpainting backbones, including SD1.5 [45], SD2.0 [47], SDXL [24], and FLUX [21].
Our main contributions can be summarized as follows:
  • We propose DreamMix, a diffusion-based method that enables precise object insertion and attribute-level editing through synergistic training-inference decoupling of identity and attributes.
  • We propose an Attribute Decoupling Mechanism (ADM) to mitigate overfitting by automatically synthesizing attribute-augmented training pairs via VLMs.
  • We propose a Textual Attribute Substitution (TAS) module that suppresses source attributes through orthogonal text embedding decomposition, ensuring target instructions dominate generation.
2 Related Work
Image Customization. Recent advances in generative models have significantly improved image customization techniques, allowing for more nuanced and flexible control over the content of generated images [62, 55, 48, 60, 17, 56]. Recently, subject-driven image customization methods [8, 44, 7, 19, 54, 51] have garnered increasing attention : given a handful of reference images, they adapt a diffusion model to faithfully reproduce a specific subject across diverse scenes and poses with high visual fidelity. For instance, Textual Inversion [8] tackles personalization by learning a dedicated text embedding from a small set of images depicting the same subject, providing a lightweight and data-efficient solution but with limited capacity for capturing complex visual details. DreamBooth [44] binds a unique identifier with the target concept and finetunes both the U-Net [42] and text encoder [43] on subject-specific images, yielding strong identity preservation and effectively preventing overfitting. OminiControl[51] concatenates conditional images and background images, and trains Diffusion Transformer model[20] based on its self-built large-scale dataset. However, these methods primarily focus on regenerating entire images but struggle with situations where users want to edit specific local regions of an existing image.
Image Inpainting. Image inpainting[32, 27, 38] involves reconstructing missing or modified regions of an image. Powered by text-to-image diffusion models, recent works [31, 2, 1, 64, 57, 28, 63, 22, 21] have shown significant improvements in repainting the local regions of existing images with text guidance. For example, Blend Diffusion [2] combines local text-guided latents with noisy versions of the input image, and a follow-up diffusion step can remap the mix to enhance coherence. SD-Inpainting [40] expands the pretrained diffusion models with a concatenation of noise, masked image, and mask as input to produce visually consistent inpainting results. PowerPaint [64] enhances the model’s ability to support various inpainting tasks by designing learnable task prompts for text-guided object restoration, context-aware image restoration, and object removal, etc. Flux-fill[21] uses Diffusion Transformer as the base model and concatenates the background image latent and mask along the channel dimension to achieve high-quality inpainting. However, despite their effectiveness in text-driven image inpainting, these methods often encounter limitations in compositing user-specific objects into the desired regions of a scene.
Subject-driven Inpainting. Recent methods have started focusing on combining subject-driven image customization with image inpainting to form a task named subject-driven inpainting [5, 26, 37, 58, 4]. To address this task, some methods [58, 5, 4] utilize an image-guided framework that retrains diffusion models using subject images as templates. However, the absence of original text guidance limits these methods to modify object attributes effectively. This motivates some approaches to develop an image-text-guided architecture that integrates these elements into latent attention layers [37, 58, 26]. Nevertheless, most methods [37, 58] rely on a large-scale training protocol that involves large-scale text-image data construction and a lengthy retraining process. While TIGIC [26] is faster, its tuning-free protocol often struggles to provide consistent editing results. In this paper, we propose DreamMix, which introduces a disentangled inpainting framework and an attribute decoupling mechanism to achieve data-efficient, editing-enhanced subject-driven image inpainting.
3 Method
Given a background image $\boldsymbol{x}$, a binary mask $\boldsymbol{m}$, and a text prompt $\boldsymbol{p}$, DreamMix aims to inpaint the local region of $\boldsymbol{x}$ specified by $\boldsymbol{m}$ to produce an image $\hat{\boldsymbol{x}}$. The refilled areas should not only contain the subject in the reference image(s) $\boldsymbol{x}_{s}$ but also align with the descriptions provided in the text prompt. The overall pipeline of DreamMix is illustrated in Figure 3. It contains three key modules: an attribute decoupling mechanism, a textual attribute substitution module, and a disentangled inpainting framework. We describe each module in detail in the following sections.
3.1 Preliminaries
Text-to-Image Diffusion Models. Diffusion models [35, 40] generate samples from random noise by learning to estimate the distribution in the reverse process. Specifically, given a random noise $\boldsymbol{\epsilon}\sim\mathcal{N}({0},\mathbf{I})$ and text prompt $\boldsymbol{p}$, text-to-image diffusion models use a noise predictor $\boldsymbol{\epsilon}_{\theta}$, often U-Net or Transformer model, to produce a denoised image. Given a dataset of $\mathcal{D}=\{\boldsymbol{x},\boldsymbol{p}\}$, the entire model is trained using the following objective:
@@ 3.1 Preliminaries
Figure 4: Pipeline of Attribute Decoupling Mechanism (ADM). We obtain the attribute word list using a VLM agent [34], construct regularization data in diverse textual formats, and get regularization image with a diffusion model.
@@ 3.2 Attribute Decoupling Mechanism
Few-shot personalization techniques have gained significant attention for their ability to balance data efficiency with generalization capability. However, limitations remain within the few-shot training protocol. On one hand, the textual descriptions used during finetuning are often overly simplified, providing only basic information about the subject’s appearance. For example, in DreamBooth [44], an image is typically described as “a [classname]”. On the other hand, these methods compress all the object’s details (e.g., color, texture, and shape) into a single identity token [sks]. Under this approach, the features of the subject are confined to the identity token [sks], making it difficult to modify the attributes of the target. Specifically, for a given new attribute, it could be either covered by original features, leaving the subject unchanged, or it may override the original features, causing compromised identity.
@@
In response, we propose an Attribute Decoupling Mechanism (ADM) to address the data scarcity and subject overfitting issues. To circumvent the heavy workload of manual annotation, we leverage advanced Vision-Language Models [34] to establish an automatic data reconstruction pipeline. The overall framework of ADM is illustrated in Figure 4. Specifically, given a set of subject image-text pairs $\mathcal{D}_{s}=\{\boldsymbol{x}_{s},\boldsymbol{p}_{s}\}$, we feed the subject images $\boldsymbol{x}_{s}$ into a VLM to generate an attribute dictionary that enumerates all attribute words associated with the given subject. Then, we ask the VLM to randomly combine the detected attributes with the original subject to form a series of text prompts with detailed descriptions such as “a brown clay teapot”. Finally, we generate new images based on the reconstructed text prompts to form a regularization benchmark $\mathcal{D}_{r}=\{\boldsymbol{x}_{r},\boldsymbol{p}_{r}\}$. Note that different from the subject benchmark $\mathcal{D}_{s}$, the identity toke sks is removed in the text prompts of regularization benchmark $\mathcal{D}_{r}$.
@@
Combining source data with regularization data to finetune the pretrained inpainting models can help reduce overfitting, yet is non-trivial. DreamBooth finetunes the model parameters beyond just identity tokens, which can lead to attribute words mixing up details. For example, in “a brown clay teapot”, the color word might inadvertently bind some texture details, making it difficult to change the texture when transitioning to “a brown glass teapot”, or influence the precise of texture features when switching to “a red clay teapot”. We refer to this phenomenon as “Concept Infusion”. Similarly, SID [18] notes that adding detailed descriptions can disrupt the object’s identity, aligning with this Concept Infusion issue.
@@
To mitigate this issue, we utilize the regularization data $\mathcal{D}_{r}$ from ADM for model finetuning and redesign the data reconstruction loss to fit the inpainting task. Specifically, we employ user masks to enforce the inpainting models to fill specific regions. And we propose a loss weight reallocation strategy on Eq. 3 to reduce the influence of background on feature learning:
[…]
@@
where $\tau_{1}=1.5$ and $\tau_{2}=0.7$ to emphasize the fill-in regions while suppressing the background area. $\hat{\boldsymbol{x}}=\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x};\boldsymbol{p})$ represents the pretrained inpainting model that generates the denosied output $\hat{\boldsymbol{x}}$ from a background image $\boldsymbol{x}$, a text prompt $\boldsymbol{p}$, and a binary mask $\boldsymbol{m}$. As illustrated in Figure 3 (a), we combine both subject data $\mathcal{D}_{s}$ and regularization data $\mathcal{D}_{r}$ with this reweight loss to form an overall loss: